# Mass and energy cascade in collisionless dark matter flow and relevant constraints on the dark matter particle massA.6

### v3

@@ A.4 The probability distribution of waiting time
A.5 Heterogeneous diffusion model and halo mass functions
The heterogeneous diffusion with spatially dependent diffusivity plays an important role in many physical problems. Examples are the mass transport in porous, inhomogeneous media and plasmas. The transport in these examples involves the waiting time that is explicitly dependent on the position. Particularly, the power-law dependence of the diffusivity is natural for many systems exhibiting self-similarity, for example, the disordered materials, the diffusion on fractals, and the mass cascade of self-gravitating collisionless dark matter flow (SG-CFD) in this work. Here, a heterogeneous diffusion model can be established for inverse mass cascade with waiting time explicitly dependent on the halo mass ($\tau_{g}\sim m_{h}^{-\lambda}$). First, the group mass $m_{g}$ for halo groups reads
@@ A.5 Heterogeneous diffusion model and halo mass functions
where $M_{h}$ is the total mass in halo sub-system, $f_{M}\left(m_{h},a\right)$ is the probability distribution with respect to the haloes mass $m_{h}$ (mass function), and $m_{p}$ is the mass resolution (particle mass). The time variation of $m_{g}$ has two contributions from Eq. (170),
@@
[…]
The governing equation for the mass function $f_{M}\left(m_{h},a\right)$ can be eventually found with substitution of Eq. (170) into Eq. (172),
@@
The solution to Eq. (175) is a stretched Gaussian function that has an exponential cut-off for large halo mass $m_{h}$ and a power-law behavior for small $m_{h}$,
@@
The mean square displacement of $m_{h}$ in mass space can be defined
[…]
where $m_{h}^{*}$ is the characteristic mass scale and $\gamma_{0}$ is a proportional constant. With the exponent of ${1/\left(1-\lambda\right)}\geq 1$ in Eq. (177), diffusion in mass space is of a super-diffusion nature. Solution of $f_{M}\left(m_{h},a\right)$ (Eq. (176)) can be further expressed in terms of $m_{h}^{*}$,
@@
Since $m_{h}^{*}$ is related to the mean square displacement of halo mass, diffusivity in halo mass space can be expressed in terms of $m_{h}^{*}$,
@@
For a constant diffusivity $D_{m0}$ with respect to a, we also would expect $\tau_{0}=1$, i.e. the fundamental frequency $f_{0}\left(a\right)\sim a^{-1}$.
@@
The mass flux and transfer functions can be obtained from definitions (Eqs. (122), (125)) in Section A.3. The mass transfer function reads
@@
Here $T_{m}\left(m_{h},a\right)\propto m_{g}\left({m_{h}/m_{p}}\right)^{2-2\lambda}$ is a typical feature of heterogeneous diffusion. The mass flux function is finally given by,
@@
Figure 22 plots the variation of the halo mass function $f_{M}$ (Eq. (178)), mass transfer function $-\Pi_{m}$ (Eq. (186)) and mass flux function $T_{m}$ (Eq. (185)) with halo size $n_{p}$ for a given set of parameters at the present time $t_{0}$. Two distinct ranges can be clearly identified from Fig. 22: the mass propagation range with a constant mass flux $\varepsilon_{m}=\Pi_{m}$ for $m_{h}<m_{h}^{*}$ and mass deposition range for $m_{h}>m_{h}^{*}$.
A.6 Existing halo mass functions
The abundance of haloes, i.e., halo mass function $f_{M}$, is one of the most fundamental quantities for analytically or semi-analytically modeling of structure formation and evolution. The Press-Schechter (PS) formalism is one of the first landmarks on the halo mass function (Press and Schechter, 1974; Bower, 1991) that can be used to predict the shape and evolution of the halo mass distribution,
@@ A.6 Existing halo mass functions
where $\delta_{c}\left(a\right)\sim a^{-1}$ is the critical density that has to be determined from a spherical collapse model or a two-body collapse model (Xu, 2021). Here $\sigma_{\delta}^{2}\left(m_{h}\right)$ is the variance of the initial density fluctuation when smoothed with a tophat filter of size $R=\left({3m_{h}/4\pi\bar{\rho}_{0}}\right)^{{1/3}}$. Here $\bar{\rho}_{0}$ is the physical background density at the current epoch $a=1$. The term $\sigma_{v}^{2}$ is the halo virial velocity dispersion. The second equality in Eq. (188) comes from the linear theory prediction of $\sigma_{\delta}^{2}\left(m_{h}\right)\sim m_{h}^{-\left(1+{n_{ps}/3}\right)}$ for a power-law power spectrum with an effective index of $n_{ps}$. The third equality in Eq. (188) comes from the virial theorem for haloes of mass $m_{h}$. Here $\sigma_{v}^{2}\left(m_{h}\right)\propto{Gm_{h}/r_{h}^{{}^{-n}}}\propto m_{h}^{{1+n/3}}$. The parameter $n$ is the exponent of gravitational potential $V_{p}\left(r\right)\sim r^{n}$. Since $\delta_{c}\left(a\right)\sim a^{-1}$ from the spherical collapse model, linear theory predicts that $\sigma_{v}^{2}\left(m_{h},a\right)\sim a^{-1}m_{h}^{1+{n_{ps}/3}}$, $\sigma_{v}^{2}\left(m_{h}^{*}\right)\sim a$ , $m_{h}^{*}\sim a^{{6/\left(3+n_{ps}\right)}}$ and $v\sim a^{-2}m_{h}^{1+{n_{ps}/3}}$.
@@
Further improvement was achieved by extending the PS formalism with the elliptical collapse model (Sheth et al., 2001; Sheth and Tormen, 1999). The modified PS model (ST model, hereafter ST) reads:
@@
The best-fitted parameters from large-scale N-body simulations are $q=0.75$ and $p=0.3$ (Sheth and Tormen, 2002). Many other forms of empirical mass functions were also proposed by fitting to the high-resolution simulation data. For example, a universal JK mass function was proposed to cover a wide range of simulation data with different cosmologies and redshifts (Jenkins et al., 2001),
@@
Now, let us return to our halo mass function from inverse mass cascade (Eq. (178)), which does not rely on any particular collapse model (spherical or elliptical). If we introduce $v=\left({m_{h}/m_{h}^{*}}\right)^{2-2\lambda}$, the halo mass function Eq. (178) can be simplified to
@@
Clearly, Eq. (178) reduces to the Press-Schechter (PS) mass function if $\lambda={\left(3-n_{ps}\right)/6}$ and $\eta_{0}={1/2}$ (See Eqs. (188)). However, it should be noted that Eq. (178) is more general and the parameter $\eta_{0}$ is related to the parameter $\beta_{0}$ (Eq. (180)), where $\beta_{0}$ is the prefactor for power-law scaling in Eq. (131). The halo geometry exponent $\lambda$ has a fundamental connection to the effective index of power spectrum $n_{ps}$. In principle, the halo geometry exponent $\lambda$ should be smaller than one such that $n_{ps}>-3$.
A universal halo mass function like Eq. (192) is clearly a manifestation of a statistically steady state involving mass and energy cascade with scale-independent rates. All these results provide insights into a fundamental question: how does the non-equilibrium system maximize its entropy and approach the limiting equilibrium via a cascade process? Two typical examples are, of course, the hydrodynamic turbulence and SG-CFD (self-gravitating collisionless dark matter flow). Both examples exhibit an intermediate and statistically steady state where a direct energy (or inverse mass) cascade process is well established.

### v4

@@ C.4 The probability distribution of waiting time
C.5 Heterogeneous diffusion model and halo mass functions
The heterogeneous diffusion with spatially dependent diffusivity plays an important role in many physical problems. Examples are mass transport in porous, inhomogeneous media, and plasmas. The transport in these examples involves the waiting time that is explicitly dependent on the position. Particularly, the power-law dependence of the diffusivity is natural for many systems exhibiting self-similarity, for example, the disordered materials, the diffusion on fractals, and the mass cascade of self-gravitating collisionless dark matter flow (SG-CFD) in this work. Here, a heterogeneous diffusion model can be established for inverse mass cascade with waiting time explicitly dependent on the halo mass ($\tau_{g}\sim m_{h}^{-\lambda}$). First, the group mass $m_{g}$ for the halo groups reads
@@ C.5 Heterogeneous diffusion model and halo mass functions
where $M_{h}$ is the total mass in the halo sub-system, $f_{M}\left(m_{h},a\right)$ is the probability distribution with respect to the mass of the halo $m_{h}$ (mass function) and $m_{p}$ is the mass resolution (particle mass). The time variation of $m_{g}$ has two contributions from Eq. (186),
@@
$\frac{1}{m_{p}}\frac{\partial}{\partial a}\int_{0}^{\infty}m_{g}\left(m_{h},a\right)dm_{h}\equiv\frac{\varepsilon_{m}}{Ha}=-\frac{\partial M_{h}}{\partial a}.$ (190)
The governing equation for the mass function $f_{M}\left(m_{h},a\right)$ can eventually be found with the substitution of Eq. (186) into Eq. (188),
@@
The solution to Eq. (191) is a stretched Gaussian function that has an exponential cut-off for a large halo mass $m_{h}$ and a power-law behavior for small $m_{h}$,
@@
The mean square displacement of $m_{h}$ in the mass space can be defined
$\begin{split}\left\langle m_{h}^{2}\right\rangle&=\int_{0}^{\infty}f_{M}\left(m_{h},a\right)m_{h}^{2}dm_{h}\\ &=\frac{1}{\sqrt{\pi}}\Gamma\left(\frac{3-\lambda}{2-2\lambda}\right)\left(2-2\lambda\right)^{\frac{2}{1-\lambda}}\left(D_{m0}a\right)^{\frac{1}{1-\lambda}}\equiv\gamma_{0}m_{h}^{*2}.\end{split}$ (193)
where $m_{h}^{*}$ is the characteristic mass scale and $\gamma_{0}$ is a proportional constant. With the exponent of ${1/\left(1-\lambda\right)}\geq 1$ in Eq. (193), diffusion in mass space is of a super-diffusion nature. The solution of $f_{M}\left(m_{h},a\right)$ (Eq. (192)) can be expressed in terms of $m_{h}^{*}$,
@@
Since $m_{h}^{*}$ is related to the mean square displacement of the halo mass, the diffusivity in the halo mass space can be expressed in terms of $m_{h}^{*}$,
@@
For a constant diffusivity $D_{m0}$ with respect to a, we would also expect $\tau_{0}=1$, i.e. the fundamental frequency $f_{0}\left(a\right)\sim a^{-1}$.
@@
The mass flux and transfer functions can be obtained from the definitions (Eqs. (138), (141)) in Section C.3. The mass transfer function reads
@@
Here $T_{m}\left(m_{h},a\right)\propto m_{g}\left({m_{h}/m_{p}}\right)^{2-2\lambda}$ is a typical feature of heterogeneous diffusion. The mass flux function is finally given by
@@
Figure 24 plots the variation of the halo mass function $f_{M}$ (Eq. (194)), the mass transfer function $\Pi_{m}$ (Eq. (202)) and the mass flux function $T_{m}$ (Eq. (201)) with the halo size $n_{p}$ for a given set of parameters at present $t_{0}$. Two distinct ranges can be clearly identified from Fig. 24: the mass propagation range with a constant mass flux $\varepsilon_{m}=\Pi_{m}$ for $m_{h}<m_{h}^{*}$ and the mass deposition range for $m_{h}>m_{h}^{*}$.
C.6 Existing halo mass functions
The abundance of haloes, i.e., halo mass function $f_{M}$, is one of the most fundamental quantities for analytically or semi-analytically modeling of structure formation and evolution. The Press-Schechter (PS) formalism is one of the first landmarks on the halo mass function (Press:1974-Formation-of-Galaxies-and-Clus; Bower:1991-The-Evolution-of-Groups-of-Gal) that can be used to predict the shape and evolution of the halo mass distribution,
@@ C.6 Existing halo mass functions
where $\delta_{c}\left(a\right)\sim a^{-1}$ is the critical density that has to be determined from a spherical collapse model or a two-body collapse model (Xu:2021-A-non-radial-two-body-collapse). Here $\sigma_{\delta}^{2}\left(m_{h}\right)$ is the variance of the initial density fluctuation when smoothed with a tophat filter of size $R=\left({3m_{h}/4\pi\bar{\rho}_{0}}\right)^{{1/3}}$. Here $\bar{\rho}_{0}$ is the physical background density at the current epoch $a=1$. The term $\sigma_{v}^{2}$ is the halo virial velocity dispersion. The second equality in Eq. (204) comes from the linear theory prediction of $\sigma_{\delta}^{2}\left(m_{h}\right)\sim m_{h}^{-\left(1+{n_{ps}/3}\right)}$ for a power-law power spectrum with an effective index of $n_{ps}$. The third equality in Eq. (204) comes from the virial theorem for haloes of mass $m_{h}$. Here $\sigma_{v}^{2}\left(m_{h}\right)\propto{Gm_{h}/r_{h}^{{}^{-n}}}\propto m_{h}^{{1+n/3}}$. The parameter $n$ is the exponent of gravitational potential $V_{p}\left(r\right)\sim r^{n}$. Since $\delta_{c}\left(a\right)\sim a^{-1}$ from the spherical collapse model, linear theory predicts that $\sigma_{v}^{2}\left(m_{h},a\right)\sim a^{-1}m_{h}^{1+{n_{ps}/3}}$, $\sigma_{v}^{2}\left(m_{h}^{*}\right)\sim a$ , $m_{h}^{*}\sim a^{{6/\left(3+n_{ps}\right)}}$ and $v\sim a^{-2}m_{h}^{1+{n_{ps}/3}}$.
@@
Further improvement was achieved by extending the PS formalism with the elliptical collapse model (Sheth:2001-Ellipsoidal-collapse-and-an-im; Sheth:1999-Large-scale-bias-and-the-peak-). The modified PS model (ST model, hereafter ST) is:
@@
The best-fitted parameters from large-scale N-body simulations are $q=0.75$ and $p=0.3$ (Sheth:2002-An-excursion-set-model-of-hier). Many other forms of empirical mass functions were also proposed by fitting to the high-resolution simulation data. For example, a universal JK mass function was proposed to cover a wide range of simulation data with different cosmologies and redshifts (Jenkins:2001-The-mass-function-of-dark-matt),
@@
Now, let us return to our halo mass function from the inverse mass cascade (Eq. (194)), which does not rely on any particular collapse model (spherical or elliptical). If we introduce $v=\left({m_{h}/m_{h}^{*}}\right)^{2-2\lambda}$, the halo mass function of Eq. (194) can be simplified to
@@
Clearly, Eq. (194) reduces to the Press-Schechter (PS) mass function if $\lambda={\left(3-n_{ps}\right)/6}$ and $\eta_{0}={1/2}$ (See Eq. (204)). However, it should be noted that Eq. (194) is more general and the parameter $\eta_{0}$ is related to the parameter $\beta_{0}$ (Eq. (196)), where $\beta_{0}$ is the prefactor for power-law scaling in Eq. (147). The halo geometry exponent $\lambda$ has a fundamental connection to the effective index of the power spectrum $n_{ps}$. In principle, the halo geometry exponent $\lambda$ should be smaller than one, such that $n_{ps}>-3$.
A universal halo mass function like Eq. (208) is clearly a manifestation of a statistically steady state involving a mass and energy cascade with scale-independent rates. All these results provide insights into a fundamental question: how does the non-equilibrium system maximize its entropy and approach the limiting equilibrium via a cascade process? Two typical examples are, of course, the hydrodynamic turbulence and SG-CFD (self-gravitating collisionless dark matter flow). Both examples exhibit an intermediate and statistically steady state, where a direct energy (or inverse mass) cascade process is well established.